On five exams, Alex has scored 84, 86, 71, 91, and 98 points out of 100. He wants to earn at least a B in the course, which corresponds to an overall average grade of 82 or better. Out of 100 points, what is the minimum number of points Alex must earn on the final exam to achieve this goal?

Answer by checkley77(12844)

(84+86+71+91+98+x)/6>=82
(430+x)/6>=82
430+x>=6*82
430+x>=492
x>=492-430
x>=62 is the minimum grade on the final test to maintain an 82 average or better.